# SMART: Self-Aware Agent for Tool Overuse Mitigation

### v1

@@ A.1 Agent Experiment Details
For LM-driven agent systems, we first prompt GPT-4o with all the questions from the GSM8K test set without using any tools. We then filter out only the questions that GPT-4o can correctly answer through pure text-based reasoning. From this refined dataset, we randomly sample 50 questions to evaluate AgentGPT and XAgent’s performance. Surprisingly, despite the core model being capable of solving all sampled questions without external tools, it still heavily relies on tools during reasoning, leading to tool overuse. Figure 3 presents a case study illustrating a specific instance of tool overuse by both agents.

### v2

@@ 7 Conclusion
Acknowledgment
This research is based upon work supported DARPA ITM Program No. FA8650-23-C-7316, DARPA ECOLE Program No. #HR00112390060 and the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Agent Experiment Details
For LM-driven agent systems, we first prompt GPT-4o with all the questions from the GSM8K test set without using any tools. We then filter out only the questions that GPT-4o can correctly answer through pure text-based reasoning. From this refined dataset, we randomly sample 50 questions to evaluate AgentGPT and XAgent’s performance. Surprisingly, despite the core model being capable of solving all sampled questions without external tools, it still heavily relies on tools during reasoning, leading to tool overuse.
@@ C.5 Confidence Validation
Appendix D Additional Evaluation and Analysis
To address concerns regarding dataset and model selection, we conducted additional experiments targeting two key areas: (1) the applicability of SMARTAgent on more complex reasoning benchmarks beyond GSM8K, and (2) the behavior of o1-like models with respect to tool use.
D.1 Evaluation on Advanced Reasoning Dataset
To assess SMARTAgent’s performance on more challenging reasoning tasks, we evaluated it on the AMC’23 benchmark (of America, MAA), a dataset known for its mathematical complexity and nuanced problem-solving requirements.
We tested two base models: Llama-3.1-8B-Instruct and Mistral-Nemo-Instruct. We compared SMARTAgent against a baseline tool prompting strategy where the tool is always made available without dynamic control.
Method Llama-3.1-8B-Instruct Mistral-Nemo-Instruct
Tool Prompt Baseline 12.50 15.00
SMARTAgent 17.50 20.00
Table 9: Accuracy (%) on AMC (2023) benchmark.
Table 9 shows that SMARTAgent outperforms the baseline across both model backbones, highlighting its ability to handle complex tasks with improved reasoning-tool use balance.
D.2 Behavior of o1-like Models
To further investigate different model’s tool use behavior, we conducted experiments on Deepseek-R1-Distilled variants of Llama and Qwen. Surprisingly, these models exhibited a tendency toward tool underuse, contrary to the overuse issue our paradigm primarily targets.
Model Time AMC (2023)
Deepseek-R1-Distilled-Llama 60.00 12.50
Deepseek-R1-Distilled-Qwen 72.00 37.50
Table 10: Tool Underuse Rate (%) on Time and AMC tasks.
We define tool underuse as the rate at which a model fails to invoke a tool in scenarios where tool use would be expected. As shown in Table 10, both distilled models significantly underutilize tools, which we attribute to potential overfitting to parametric reasoning, a phenomenon aligned with “overthinking” reported in prior literatures.
D.3 SMARTAgent Adaptation for Distilled Models
To further test the adaptability of SMARTAgent, we fine-tuned these distilled models using our SMART paradigm and evaluated them on a time-domain QA benchmark.
Method Distilled-Llama Distilled-Qwen
Base Model Reasoning Prompt 30.00 12.00
Base Model Tool Prompt 36.00 26.00
SMARTAgent 40.00 52.00
Table 11: Accuracy (%) on Time-domain QA with Distilled Models.
Table 11 shows that SMARTAgent enhances both models’ performance, demonstrating its effectiveness not only in mitigating overuse but also in addressing underuse by promoting strategic tool engagement. These findings reinforce SMART’s broader applicability across diverse reasoning paradigms and model types.